• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name eschulte

    Log entry time 00:02:23 on August 05, 2006

    Entry number 177476

    keyword=Swing Shift Log

    16:00 Start shift with no beam: injector/Hall A laser/Hall A arc
    problems. The half-wave plate is out.

    16:10 Thunder storm starts, possibly causing machine problems.

    16:20 MCC calls to let us know that the PSS is having a problem
    (possibly storm related) prohibiting the delivery of beam to the halls.
    They have called their expert to help them reset the system.

    16:41 Still having problems delivering beam, not sure why yet, but we're
    having nothing but FSD's. Also, the BCM strip chart's limits seem to be
    messed up for Hall A's current, and I can't seem to reset them. The
    numerical display, however, seems to be OK.

    16:45 Got the BCM strip chart with usable limits. Needed to reset the
    limits & then restart the window.

    16:47 It looks like we're getting close to getting decent beam.

    17:16 Hall A laser is disfunctional according to MCC. They have called
    in the expert! "A couple of hours", said MCC.

    17:47 Brad requested, and we did, a pedistal run for each arm.

    18:57 MCC called to let us know that they are going to send some tune
    beam to see if they have figured out the problem.

    19:01 Requested 4uA from MCC & changed target from empyt to 15 cm LH2.

    19:08 MCC says they are only sending about 5uA while we were seeing
    about 7uA. They are going to call Brad to find out if any
    coefficients/calibrations on the BCM's were adjusted.

    19:21 Checked the rates on the 15cm LH2 target for the beam current that
    MCC is sending (4uA from MCC, 7-9 uA read on our BCM). The rates are
    about 2kHz, consistent with 4uA beam delivery on this target. This has
    been noted in the comments section of the run-start files. Beam is
    getting trippy with attendant beam position fluctuations with FFB
    switched on, according to the telemetry. Effected run 2134.

    19:42 Beam positions are closer to acceptable. They still deviate to
    about 0.3mm, but are closer, more often <0.2 mm than for the previous
    run. The FFB has been having trouble (keeps switching to standby).

    20:00 Jack called, he had checked the HRS-R values and noticed that they
    were no longer changing. We rebooted the IOC & it came back, without
    having to reset any currents, just the momentum setpoint to get rid of
    the alarms. FFB still having some trouble staying locked.

    20:12 Beam positions have been very unstable for the last hour or so. I
    have called MCC to see if they can figure out if it's the Fast Feedback
    again.

    20:21 MCC is switching to tune beam to try to fix the beam position
    stability problems we've been seeing.

    20:46 Hall C is going to do a IA; PITA; RHWP scan for about the next
    hour which may cause large charge asymetries.

    21:14 MCC called to see what we thought about the beam positions, they
    have been drifting too much, and I though they were still steering, but
    maybe they weren't...

    21:23 OK, so we've given it an hour, and the beam positions are a little
    better. MCC has been working on the position problem, but for at least
    part of this hour it was unclear whether they were working on it or not,
    as Hall C is running some tests and they didn't call us back after the
    tune beam. For my part since the beam positions were still changing
    ALOT, I though they were still working on it. I think we had a small
    communication faliure there...

    Anyway, we're trying to take data again. The positions may still jump
    around, though. New run 2138 (HRS-L).

    21:28 Run 2138 effected by MCC sending more tune beam to try to fix up
    the beam positions.

    21:55 Since MCC was sending us tune beam again to try to fix the beam
    position problems we've been having, we decided to pack it in for KIN#
    10a and set up for KIN #3 in the event we get the beam back.

    The spectrometers are now set up and we will move the target when we get
    beam back. This has been a very frustrating shift.

    22:28 Start receiving, breifly 4uA cw, with bad beam positions.

    22:51 Going to try to take data at the new kinematics setting (KIN #3).
    Put in the carbon optics target and requested the beam back at 4uA. We
    won't have any fast feedback for at least this run.

    OK, so we called MCC to get an update and they said that they've gotten
    rid of the beam stability problems, but the fast feedback is not working.
    They're trying to get it up and running for Hall C. We can take the
    beam with no FFB after they reset the beam positions.

    We will try to take some data despite all the odds against us.

    23:15 Beam positions are still unacceptable for data taking BPMA-y=1.6,
    BPMB-y=2

    23:17 Re-adjusted loop 3 JT to compensate for an increase in flow (up
    to about 12g/s) from the ESR. Reduced HPH load to about 40 W.

    23:25 Noticed that MCC seems to have Hall C's fast feedback back on, at
    least somewhat.

    23:49 Finally, we have acceptable beam positions for data taking. so
    we're going to try to get some carbon optics target. We seem to be
    having communication problems with MCC. We asked twice to have the beam
    positions reset in the last hour, and they were just reset.

    23:54 Requested increase in current to 20uA on carbon optics target.